OpenStudy (truwhovian11):

To the nearest hundredth, how many cm are in 52 in? (1 in = 2.54 cm) A) 127.40 cm B) 132.08 cm C) 179.40 cm D) 184.08 cm

OpenStudy (truwhovian11):

@CursedWizard @briannaontimeforreal @BrownieCat

OpenStudy (phi):

to do this, you notice you have 52 in if you multiply by a number that has units cm/in the "in" cancel start with 1 in = 2.54 cm you want a ratio that has cm on top, so write the ratio \[ \frac{2.54 \ cm}{1 \ in} \] next multiply \[ 52 \ \cancel{in} \cdot \frac{2.54 \ cm}{1 \ \cancel{in}} \]

OpenStudy (briannaontimeforreal):

what do you think

OpenStudy (truwhovian11):

what @phi @briannaontimeforreal i think its a or d

OpenStudy (phi):

do you know how to multiply ?

OpenStudy (briannaontimeforreal):

nun of those are

OpenStudy (truwhovian11):

yes @phi

OpenStudy (briannaontimeforreal):

right it c or b

OpenStudy (truwhovian11):

is it C

OpenStudy (briannaontimeforreal):

correct

OpenStudy (truwhovian11):

yaaaaaaaaaaaaaaaaaaayyyyyyyyyyyyyyyyyyyyyy thanks but there is still more

OpenStudy (briannaontimeforreal):

ok